40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1013     Entry time: Tue Jul 3 19:29:07 2012
Author: Sarah 
Type: DailyProgress 
Category: Seismic 
Subject: Seismic Stack 

Tara and I tried to measure the resonant frequency of the seismic stack in horizontal modes, looking for a correlation to some of the peaks in the noise budget. 

Setup/ Impulse Response:

We first removed the pressure from the table so that it was not raised at all. Then after locking the beams to the corresponding cavities, we slightly misaligned the beam in RCAV. The beam misalignment allowed us to see the entire impulse response from the pd behind RCAV when the cavity was hit with a hammer, exciting the horizontal modes. We used the oscilloscope to get the data for the impulse response.

The following plot shows several impulse responses:

Impulse_Response.png

Red: DC , 1000Hz sampling rate

Green: DC , 500Hz sampling rate

Blue: DC , 500Hz sampling rate

Purple: DC, 500Hz sampling rate

Frequency Response:

A FFT was used to convert the impulse response to frequency response. We hoped to see a frequency response peaks that corresponded to noise peaks on the noise budget for the beat measurement. The following plot shows the Frequency Responses for the above Impulse Responses:

Frequency_Response.png

For comparison, the noise budget is:

nb_allfreq.png

It is difficult to see whether the peaks correspond to anything, because they die out at the frequencies of interest for the noise budget.

Comments:

We noticed that the DC offset was changing during the impulse response measurements, and thus it is possible that we might be observing the frequency response of the rolling of the cavity as opposed to the frequency response of the seismic stacks.

Attachment 4: nb_allfreq.fig  310 kB
Attachment 5: Frequency_Response.fig  125 kB
Attachment 6: Impulse_Response.fig  41 kB
ELOG V3.1.3-